the length of the top of a table is 5 m greater than the width. The area is 104 m^2. Find the dimensionsof the tables
Answer by jorel1380(3719)

x(x+5)=104
x2+5x-104=0
(x+13)(x-8)=0
x=-13 or 8
Throwing out the negative result, we get the dimensions of the table to be 8' x 13'..